# Melbourne, VIC 3010, Australia

### v2

@@ 2.1 Topological strings and intersection numbers
where $N_{\mathfrak{g},\beta}$ are the genuine Gromov–Witten invariants in genus $\mathfrak{g}$ and degree $\beta$, cf. (2), ¯$\overline{{\cal M}}_{\mathfrak{g}}$ is the moduli space of genus $\mathfrak{g}$ Riemann surfaces and $c_{\mathfrak{g}-1}$ is the Chern class of the Hodge bundle over ¯$\overline{{\cal M}}_{\mathfrak{g}}$. From here on, we focus the discussion on the constant maps, which are the first terms in the expressions above, also denoted $N_{\mathfrak{g},0}$.
@@
$\mathrm{vol}_{X}(t):=$ $\sum_{a_{1},\dots,a_{d}}\frac{1}{d!}C_{a_{1},\dots,a_{d}}\,t^{a_{1}}\dots t^{a_{d}}$ (15)
@@ 2.2 Equivariant volumes, toric geometry, and generating functions
In the toric case, $X$ can be described as a Kähler quotient of a linear space $\mathbb{C}^{n}$ by a torus $U(1)^{r}$, such that the complex dimension of $X$ is given by $d=n-r$. The action of this torus is described in terms of a matrix of integer charges $Q_{i}^{a}$ with $i=1,\dots,n$ and $a=1,\dots,r$, corresponding to an embedding of $U(1)^{r}$ into the larger torus $U(1)^{n}$ acting diagonally on the homogeneous coordinates $z_{i}$ on $\mathbb{C}^{n}$. The associated moment map $\mu(z)=\rho_{a}Q_{i}^{a}|z_{i}|^{2}$ is the composition of the map $Q$ and the moment map $H=\rho_{i}H^{i}=\rho_{i}|z_{i}|^{2}$ for the $U(1)^{n}$-action, where $\rho_{a}$ and $\rho_{i}$ are a basis of $\mathrm{Lie}(U(1)^{r})^{\ast}\cong\mathbb{R}^{r}$ and $\mathrm{Lie}(U(1)^{n})^{\ast}\cong\mathbb{R}^{n}$, respectively.
@@
where the choice of contour depends on the chamber for the Kähler parameters $t^{a}$, also known as phase of the symplectic quotient Goldin:2003dis . Again, we can regard $\mathrm{vol}_{X}(t,\epsilon)$ as a generating function of certain geometric quantities, namely, the equivariant integrals over $X$ of all polynomial functions in the tautological classes $\phi_{a}$. An alternative way to construct a generating function of equivariant integrals is as follows.
@@
Compared to $\mathrm{vol}(t,\epsilon)$, the function $\mathbb{V}(\lambda,\epsilon)$ does not depend on all of the parameters $\epsilon_{i}$ but rather on $d=n-r$ independent linear combinations of them. This follows by observing the following.
@@ 2.2.1 $\mathfrak{g}=0$
It could be counterintuitive at first, but we keep this definition, using the triple intersection numbers, for a toric manifold of arbitrary complex dimension. We argue that this is the correct choice based on explicit comparison with M2-brane partition functions, Cassia:2024 , see also Martelli:2023oqk .
@@ 2.2.2 $\mathfrak{g}=1$
where $x_{i}=\epsilon_{i}+Q_{i}^{a}\phi_{a}$ and $d=n-r$ is the complex dimension of the quotient. Because of equivariance again, this is not a polynomial of fixed degree in $t^{a}$’s, but a power series.
@@ 2.2.3 $\mathfrak{g}\geq 2$
¯$\overline{{\cal M}}_{\mathfrak{g}}(X,0)\cong X\times\overline{{\cal M}}_{\mathfrak{g}}$ (50)
@@
is a constant also independent of $\epsilon_{i}$’s because of cohomological degree reasons. In general however, this function is homogeneous of degree $3-d$ w.r.t. a simultaneous rescaling of all the $\epsilon_{i}$’s.
@@ 2.3 Calabi–Yau cones, Kähler bases, and mesonic twist
In turn, one can regard such parameters $t^{a}$ as measuring the lengths of the compact 1-dimensional legs in the toric diagram of $X$. By definition, the toric diagram is the 1-skeleton of the intersection locus of the preimage of the moment maps with the positive orthant in $\mathbb{R}^{n}$. The CY condition on $Q_{i}^{a}$ implies that this intersection in non-compact, and very far from the origin it looks like a real $2d$-dimensional cone, $d=(n-r)$. When we set all the Kähler moduli to zero, we then expect that $X$ becomes a cone where all the non-trivial homology 2-cycles have been shrunk to a single point coinciding with the tip of the cone. The base of this cone is a Sasakian manifold $L$, of real dimension $12d-$ 1. Conversely, we can regard $X$ as a toric resolution of the cone over $L$. Because $t=0$ is at the origin of the extended Kähler moduli space, after shrinking all homology 2-cycles in $X$, we end up with a conical singular space which no longer depends on the choice of chamber (all chambers contain the origin as a boundary point). From now on we will refer to this conical space as $X_{\circ}$ and we will use that
@@
$N_{1,0}^{\text{mes.}}(\lambda_{\text{mes.}},\epsilon)$ $=\frac{1}{24}\,\left(\epsilon_{i}\lambda_{\text{mes.}}^{i}\right)\,\sum_{i<j}\frac{\partial^{2}\mathbb{V}(\lambda,\epsilon)}{\partial\lambda^{i}\partial\lambda^{j}}\Big{|}_{\lambda=0}\,\ ,$
@@ 3.2 Spindle, $\mathbb{WP}^{1}_{a,b}$
As expected, the leading order is proportional to $t$ and it parametrizes the ordinary volume, while the equivariant upgrade adds corrections at higher orders in the Kähler modulus $t$ and the equivariant parameters ϵ±subscriptitalic-$\epsilon_{\pm}$’s. [^13] From a cohomological point of view, one should assign degree $-2$ to the Kähler moduli $t^{a}$ and degree $+2$ to the equivariant parameters $\epsilon_{i}$, so that the function $\mathrm{vol}(t,\epsilon)$ becomes homogeneous of degree $-2d$. Note that in order for the quotient to be a compact two-fold as described above, we need to impose that both charges are positive integers and that the Kähler modulus is also greater than zero. However, once the volume has been computed via equivariant localization as in (91), we can analytically continue the function $\mathrm{vol}(t,\epsilon)$ to arbitrary values of $a,b$ and $t$. In particular, we will allow $b$ to become negative while keeping $a$ positive (or equivalently the other way around). The reason for doing so has to do with the fact that, from the physics point of view, this choice still corresponds to a BPS background in string theory, as argued in Martelli:2023oqk (see also Cassia:2024 ).
@@
$v=\begin{pmatrix}-b\\ a\end{pmatrix}\ ,$ (96)
@@ 3.3 Projective space, $\mathbb{P}^{k}$
$v=\begin{pmatrix}1&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&1\\ -1&-1&\dots&-1\end{pmatrix}\ ,$ (103)
@@ 4.1 Local $\mathbb{P}^{n-1}$
Unsurprisingly, after performing the mesonic twist, i.e. blowing down the non-trivial two-cycle of volume $t$, we obtain the equivariant volume of the cone $\mathbb{C}^{n}/\mathbb{Z}_{n}$ [^14] Note that the overall numerical prefactor is ambiguous as it can be reabsorbed in the definition of the equivariant parameters, such that the first equality is conventional.
@@ 4.2 Resolved conifold
$\begin{array}[]{ccc}\epsilon_{1}=\frac{(\epsilon_{1}+\epsilon_{3})(\epsilon_{1}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,&&\epsilon_{3}=\frac{(\epsilon_{1}+\epsilon_{3})(\epsilon_{2}+\epsilon_{3})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,\\ &&\\ \epsilon_{2}=\frac{(\epsilon_{2}+\epsilon_{3})(\epsilon_{2}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,&&\epsilon_{4}=\frac{(\epsilon_{1}+\epsilon_{4})(\epsilon_{2}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,\end{array}$ (126)
@@
This is precisely the constraint discussed in (Hosseini:2019ddy, , Section 5.1) [^15] The rest of the quantities discussed here also have a direct correspondence to the ones in (Hosseini:2019ddy, , Section 5.1), with $\epsilon^{\text{here}}\leftrightarrow\Delta^{\text{there}}$ and $\nu^{\text{here}}\leftrightarrow b^{\text{there}}$. One should however take into account the additional equivariant parameter in that reference that comes from a direct product with $\mathbb{C}$, which we have not considered here but is straightforward to include. which on the field theory side can be understood as lifting up the flat directions in the matrix model, see Jafferis:2011zi . The constant maps terms in the topological string amplitudes after the mesonic twist are given by (69), under the constraint (127).
@@ 5.1 $\mathbb{C}^{d}$ fiber
where $\epsilon_{i}$ are the equivariant parameters on $\mathbb{C}^{d}$, ϵ±subscriptitalic-$\epsilon_{\pm}$ those of the base $\mathbb{WP}^{1}_{a,b}$ as in Section 3.2, and we took the poles at $\phi=-\frac{\epsilon_{+}}{a}$ and $\phi=-\frac{\epsilon_{-}}{b}$. We observe that the answer has a natural interpretation in terms of the equivariant volume of $\mathbb{C}^{d}$, (81),
@@
where we already used the definition (94), $\bar{\nu}=a\epsilon_{-}-b\epsilon_{+}$, in this case simply as short-hand since $\mathrm{vol}_{Y}$ is still understood as a function of ϵ±subscriptitalic-$\epsilon_{\pm}$. In this case, the fiber has no Kähler moduli, but its equivariant parameters $\epsilon_{i}$ get shifted by the fluxes $n_{i}$ multiplied by the equivariant parameters of the normal bundle to the fixed point in $\mathbb{WP}^{1}_{a,b}$.
@@
$\begin{array}[]{c}a\,\epsilon_{-}=\frac{\bar{\nu}\,\prod_{j}\epsilon^{-}_{j}}{\prod_{j}\epsilon^{-}_{j}-\prod_{k}\epsilon^{+}_{k}}\ ,\qquad b\,\epsilon_{+}=\frac{\bar{\nu}\,\prod_{k}\epsilon^{+}_{k}}{\prod_{j}\epsilon^{-}_{j}-\prod_{k}\epsilon^{+}_{k}}\ ,\\ \\ \epsilon_{i}=\frac{\epsilon^{+}_{i}\,\prod_{j}\epsilon^{-}_{j}-\epsilon^{-}_{i}\,\prod_{k}\epsilon^{+}_{k}}{\prod_{j}\epsilon^{-}_{j}-\prod_{k}\epsilon^{+}_{k}}\ ,\end{array}$ (171)
@@
$a=b=1\ ,\qquad\Rightarrow\qquad n=2\ ,$ (172)
@@
with ϵ±subscriptitalic-$\epsilon_{\pm}$ unconstrained. Plugging this back into the equivariant volume, we find
@@
which is precisely matching with $\mathbb{V}_{\mathbb{C}^{2}/\mathbb{Z}_{2}}(\lambda,\epsilon)$ upon the identification of $\pm$ indices with the ${1,2}$ indices of $\mathbb{C}^{2}$. The precise match between the equivariant constant maps terms (168) and (84) also follows simply from the same identification.
@@ 5.2 General and mesonic fiber
where $Q_{X}$ is the matrix of charges of the fiber and $n_{i}$ are (non-negative) integers describing the fibration.  [^17] The structure group of the fibration corresponds to the torus group of isometries of the fiber, which we identify with $U(1)^{n}$. The charges $n_{i}$ then represent the first Chern numbers of the associated principal $U(1)$-bundles. Since not all of these $U(1)$’s act effectively, the topology of the total space $Y$ only depends on $d=\dim X$ linear combinations of the $n_{i}$, as in (182). Another way to see this, is to observe that $Q_{Y}$ is only defined up to the left-action of $SL_{r+1}(\mathbb{Z})$ which can be used to shift the vector of $n_{i}$ by the rows of the matrix $Q_{X}$. In this description, $Y$ can be explicitly identified with the quotient $\mathbb{C}^{n+2}//U(1)^{r+1}$ of complex dimension $n-r+1=d+1$ ($d$ being the dimension of $X$), which is CY provided that the charges satisfy the condition
@@
in agreement with Martelli:2023oqk  [^19] Note that the charges in Martelli:2023oqk are normalizaed to carry an additional inverse factor of $(ab)$ with respect to the present conventions. In analogy to the previous subsection, we can redefine equivariant parameters on the total space of the fibration as
@@
which is trivial even before the mesonic twist on $\lambda$ due to (40). This implies that, as in the case of the constraints (71), only $n-d$ of these equations are linearly independent, and they can be used to lift the redundancy from the $n_{i}$ parameters to the $p_{\alpha}$ parameters. As mentioned, the specialization $\lambda=\lambda_{\rm mes.}$ is not strictly necessary for this procedure to work, however it does make the expression in the r.h.s. of (194) considerably simpler.
@@
6 Flux compactifications
As explained in the introduction, our topological string examples are primarily motivated by flux compactifications, particularly M2-brane constructions preserving at least ${\cal N}=2$ supersymmetry, see Gabella:2012rc and references therein. While the relevance of cones over Sasakian manifolds extends beyond M2-branes, we focus on this specific case for clarity. It is important to emphasize that our approach from here on is guided by supersymmetry, beginning from the low-energy description. Although supergravity does not capture all details, as illustrated in Figures 1 and 2, we rely on supersymmetry to fix the topology of the underlying manifold $X$. The immediate goal here is to clarify the claim that the $\lambda$ parameters in the equivariant volume are conjugate to the fluxes $N_{\rm M2}$, proposed in terms of an extremization procedure in Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Martelli:2023oqk . In Cassia:2024 , we elaborate in more detail on many of the aspects touched upon here and reformulate the relation as a full integral transform. The present section thus serves as an introduction to the companion paper, at the same time physically motivating the examples we have considered in the previous sections.

### v4

@@ 2.1 Topological strings and intersection numbers
where $N_{\mathfrak{g},\beta}$ are the genuine Gromov–Witten invariants in genus $\mathfrak{g}$ and degree $\beta$, cf. (2), $\overline{{\cal M}}_{\mathfrak{g}}$ is the moduli space of genus $\mathfrak{g}$ Riemann surfaces and $c_{\mathfrak{g}-1}$ is the Chern class of the Hodge bundle over $\overline{{\cal M}}_{\mathfrak{g}}$. From here on, we focus the discussion on the constant maps, which are the first terms in the expressions above, also denoted $N_{\mathfrak{g},0}$.
@@
$\mathrm{vol}_{X}(t)$= $\sum_{a_{1},\dots,a_{d}}\frac{1}{d!}C_{a_{1},\dots,a_{d}}\,t^{a_{1}}\dots t^{a_{d}}$ (15)
@@ 2.2 Equivariant volumes, toric geometry, and generating functions
In the toric case, $X$ can be described as a Kähler quotient of a linear space $\mathbb{C}^{n}$ by a torus $U(1)^{r}$, such that the complex dimension of $X$ is given by d=nrd=n-r. The action of this torus is described in terms of a matrix of integer charges $Q_{i}^{a}$ with $i=1,\dots,n$ and $a=1,\dots,r$, corresponding to an embedding of $U(1)^{r}$ into the larger torus $U(1)^{n}$ acting diagonally on the homogeneous coordinates $z_{i}$ on $\mathbb{C}^{n}$. The associated moment map $\mu(z)=\rho_{a}Q_{i}^{a}|z_{i}|^{2}$ is the composition of the map $Q$ and the moment map $H=\rho_{i}H^{i}=\rho_{i}|z_{i}|^{2}$ for the $U(1)^{n}$-action, where $\rho_{a}$ and $\rho_{i}$ are a basis of $\mathrm{Lie}(U(1)^{r})^{\ast}\cong\mathbb{R}^{r}$ and $\mathrm{Lie}(U(1)^{n})^{\ast}\cong\mathbb{R}^{n}$, respectively.
@@
where the choice of contour depends on the chamber for the Kähler parameters $t^{a}$, also known as phase of the symplectic quotient Goldin:2003dis . Evaluation of the residues at each of the JK-poles allows to recover the contributions of each fixed-point as in the ABBV localization formula.
Again, we can regard $\mathrm{vol}_{X}(t,\epsilon)$ as a generating function of certain geometric quantities, namely, the equivariant integrals over $X$ of all polynomial functions in the tautological classes $\phi_{a}$. An alternative way to construct a generating function of equivariant integrals is as follows.
@@
Compared to $\mathrm{vol}(t,\epsilon)$, the function $\mathbb{V}(\lambda,\epsilon)$ does not depend on all of the parameters $\epsilon_{i}$ but rather on d=nrd=n-r independent linear combinations of them. This follows by observing the following.
@@ 2.2.1 $\mathfrak{g}=0$
It could be counterintuitive at first, but we keep this definition, using the triple intersection numbers, for a toric manifold of arbitrary complex dimension. We argue that this is the correct choice based on explicit comparison with M2-brane partition functions, Cassia:2025jkr , see also Martelli:2023oqk .
@@ 2.2.2 $\mathfrak{g}=1$
where $x_{i}=\epsilon_{i}+Q_{i}^{a}\phi_{a}$ and d=nrd=n-r is the complex dimension of the quotient. Because of equivariance again, this is not a polynomial of fixed degree in $t^{a}$’s, but a power series.
@@ 2.2.3 $\mathfrak{g}\geq 2$
$\overline{{\cal M}}_{\mathfrak{g}}(X,0)\cong X\times\overline{{\cal M}}_{\mathfrak{g}}$ (50)
@@
is a constant also independent of $\epsilon_{i}$’s because of cohomological degree reasons. In general however, this function is homogeneous of degree 3d3-d w.r.t. a simultaneous rescaling of all the $\epsilon_{i}$’s.
@@ 2.3 Calabi–Yau cones, Kähler bases, and mesonic twist
In turn, one can regard such parameters $t^{a}$ as measuring the lengths of the compact 1-dimensional legs in the toric diagram of $X$. By definition, the toric diagram is the 1-skeleton of the intersection locus of the preimage of the moment maps with the positive orthant in $\mathbb{R}^{n}$. The CY condition on $Q_{i}^{a}$ implies that this intersection in non-compact, and very far from the origin it looks like a real $2d$-dimensional cone, d=(nr)d=(n-r). When we set all the Kähler moduli to zero, we then expect that $X$ becomes a cone where all the non-trivial homology 2-cycles have been shrunk to a single point coinciding with the tip of the cone. The base of this cone is a Sasakian manifold $L$, of real dimension 2d12d-1. Conversely, we can regard $X$ as a toric resolution of the cone over $L$. Because $t=0$ is at the origin of the extended Kähler moduli space, after shrinking all homology 2-cycles in $X$, we end up with a conical singular space which no longer depends on the choice of chamber (all chambers contain the origin as a boundary point). From now on we will refer to this conical space as $X_{\circ}$ and we will use that
@@
$N_{1,0}^{\text{mes.}}(\lambda_{\text{mes.}},\epsilon)$ $=\frac{1}{24}\,\left(\epsilon_{i}\lambda_{\text{mes.}}^{i}\right)\,\sum_{i<j}\frac{\partial^{2}\mathbb{V}(\lambda,\epsilon)}{\partial\lambda^{i}\partial\lambda^{j}}\Big|_{\lambda=0}\,\$,
@@ 3.2 Spindle, $\mathbb{WP}^{1}_{a,b}$
As expected, the leading order is proportional to $t$ and it parametrizes the ordinary volume, while the equivariant upgrade adds corrections at higher orders in the Kähler modulus $t$ and the equivariant parameters $\epsilon_{\pm}$’s. [^16] From a cohomological point of view, one should assign degree 2-2 to the Kähler moduli $t^{a}$ and degree +2+2 to the equivariant parameters $\epsilon_{i}$, so that the function $\mathrm{vol}(t,\epsilon)$ becomes homogeneous of degree 2d-2d. Note that in order for the quotient to be a compact two-fold as described above, we need to impose that both charges are positive integers and that the Kähler modulus is also greater than zero. However, once the volume has been computed via equivariant localization as in (91), we can analytically continue the function $\mathrm{vol}(t,\epsilon)$ to arbitrary values of $a,b$ and $t$. In particular, we will allow $b$ to become negative while keeping $a$ positive (or equivalently the other way around). The reason for doing so has to do with the fact that, from the physics point of view, this choice still corresponds to a BPS background in string theory, as argued in Martelli:2023oqk (see also Cassia:2025jkr ).
@@
$v=\begin{pmatrix}-b\\ a\end{pmatrix}\$, (96)
@@ 3.3 Projective space, $\mathbb{P}^{k}$
$v=\begin{pmatrix}1&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&1\\ -1&-1&\dots&-1\end{pmatrix}\$, (103)
@@ 4.1 Local $\mathbb{P}^{n-1}$
Unsurprisingly, after performing the mesonic twist, i.e. blowing down the non-trivial two-cycle of volume $t$, we obtain the equivariant volume of the cone $\mathbb{C}^{n}/\mathbb{Z}_{n}$:  [^17] Note that the overall numerical prefactor is ambiguous as it can be reabsorbed in the definition of the equivariant parameters, such that the first equality is conventional.
@@ 4.2 Resolved conifold
$\begin{array}[]{ccc}\epsilon_{1}=\frac{(\epsilon_{1}+\epsilon_{3})(\epsilon_{1}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,&\quad&\epsilon_{3}=\frac{(\epsilon_{1}+\epsilon_{3})(\epsilon_{2}+\epsilon_{3})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,\\ &&\\ \epsilon_{2}=\frac{(\epsilon_{2}+\epsilon_{3})(\epsilon_{2}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,&\quad&\epsilon_{4}=\frac{(\epsilon_{1}+\epsilon_{4})(\epsilon_{2}+\epsilon_{4})}{\sum_{i=1}^{4}\epsilon_{i}}\ ,\end{array}$ (126)
@@
This is precisely the constraint discussed in (Hosseini:2019ddy, , Section 5.1),  [^18] The rest of the quantities discussed here also have a direct correspondence to the ones in (Hosseini:2019ddy, , Section 5.1), with $\epsilon^{\text{here}}\leftrightarrow\Delta^{\text{there}}$ and $\nu^{\text{here}}\leftrightarrow b^{\text{there}}$. One should however take into account the additional equivariant parameter in that reference that comes from a direct product with $\mathbb{C}$, which we have not considered here but is straightforward to include. which on the field theory side can be understood as lifting up the flat directions in the matrix model, see Jafferis:2011zi . The constant maps terms in the topological string amplitudes after the mesonic twist are given by (69), under the constraint (127).
@@ 5.1 $\mathbb{C}^{d}$ fiber
where $\epsilon_{i}$ are the equivariant parameters on $\mathbb{C}^{d}$, $\epsilon_{\pm}$ those of the base $\mathbb{WP}^{1}_{a,b}$ as in Section 3.2, and we took the poles at $\phi=-\frac{\epsilon_{+}}{a}$ and $\phi=-\frac{\epsilon_{-}}{b}$. We observe that the answer has a natural interpretation in terms of the equivariant volume of $\mathbb{C}^{d}$, (81),
@@
where we already used the definition (94), $\bar{\nu}=a\epsilon_{-}-b\epsilon_{+}$, in this case simply as short-hand since $\mathrm{vol}_{Y}$ is still understood as a function of $\epsilon_{\pm}$. In this case, the fiber has no Kähler moduli, but its equivariant parameters $\epsilon_{i}$ get shifted by the fluxes $n_{i}$ multiplied by the equivariant parameters of the normal bundle to the fixed point in $\mathbb{WP}^{1}_{a,b}$.
@@
[…]
@@
a=b=1,n=2,a=b=1\ $\qquad\Rightarrow\qquad$ n=2\ , (172)
@@
with $\epsilon_{\pm}$ unconstrained. Plugging this back into the equivariant volume, we find
@@
which is precisely matching with $\mathbb{V}_{\mathbb{C}^{2}/\mathbb{Z}_{2}}(\lambda,\epsilon)$ upon the identification of $\pm$ indices with the 1,2{1,2} indices of $\mathbb{C}^{2}$. The precise match between the equivariant constant maps terms (168) and (84) also follows simply from the same identification.
@@ 5.2 General and mesonic fiber
where $Q_{X}$ is the matrix of charges of the fiber and $n_{i}$ are (non-negative) integers describing the fibration.  [^20] The structure group of the fibration corresponds to the torus group of isometries of the fiber, which we identify with $U(1)^{n}$. The charges $n_{i}$ then represent the first Chern numbers of the associated principal $U(1)$-bundles. Since not all of these $U(1)$’s act effectively, the topology of the total space $Y$ only depends on $d=\dim X$ linear combinations of the $n_{i}$, as in (182). Another way to see this, is to observe that $Q_{Y}$ is only defined up to the left-action of $SL_{r+1}(\mathbb{Z})$ which can be used to shift the vector of $n_{i}$ by the rows of the matrix $Q_{X}$. In this description, $Y$ can be explicitly identified with the quotient $\mathbb{C}^{n+2}//U(1)^{r+1}$ of complex dimension nr+1=d+1n-r+1=d+1 ($d$ being the dimension of $X$), which is CY provided that the charges satisfy the condition
@@
in agreement with Martelli:2023oqk .  [^22] Note that the charges in Martelli:2023oqk are normalizaed to carry an additional inverse factor of $(ab)$ with respect to the present conventions. In analogy to the previous subsection, we can redefine equivariant parameters on the total space of the fibration as
@@
which is trivial even before the mesonic twist on $\lambda$ due to (40). This implies that, as in the case of the constraints (71), only ndn-d of these equations are linearly independent, and they can be used to lift the redundancy from the $n_{i}$ parameters to the $p_{\alpha}$ parameters. As mentioned, the specialization $\lambda=\lambda_{\rm mes.}$ is not strictly necessary for this procedure to work, however it does make the expression in the r.h.s. of (194) considerably simpler.
@@
6 M2-brane backgrounds with flux
As explained in the introduction, our topological string examples are primarily motivated by charged brane configurations, particularly M2-brane constructions preserving at least ${\cal N}=2$ supersymmetry, see Gabella:2012rc and references therein. While the relevance of cones over Sasakian manifolds extends beyond M2-branes, we focus on this specific case for clarity. It is important to emphasize that our approach from here on is guided by supersymmetry, beginning from the low-energy description. Although supergravity does not capture all details, as illustrated in Figures 1 and 2, we rely on supersymmetry to fix the topology of the underlying manifold $X$. The immediate goal here is to clarify the claim that the $\lambda$ parameters in the equivariant volume are conjugate to the fluxes $N_{\rm M2}$, proposed in terms of an extremization procedure in Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Martelli:2023oqk . The main results of these references in practice illustrate the validity of equations (7) and (9) at leading order in the M2-brane charge $N$. In Cassia:2025jkr , we elaborate in more detail on many of the aspects touched upon here and reformulate the relation as a full integral transform. The present section thus serves as an introduction to the companion paper, while also physically motivating the examples we have considered in the previous sections.